The amount of flux in a round magnetic rod measured at 0.013 weber.If the diameter of the material is 12 cm, calculate the flux density?
1
Expert's answer
2020-10-09T07:11:47-0400

Solution

Given data

Flux "(\Phi) =0.013" Webers

Diameter=12cm

Radius(r) =6cm=0.06m

Area(A) ="\pi r^2 =3.14\times0.06\times0.06"

=0.0113 "m^2"

Flux density can be expressed as

"B=\frac{\Phi}{A}=\frac{0.013}{0.0113}=1.15Tesla"

Therefore flux density is 1.15 T.